6. Determine the minimum sample size required to have a margin of error of 2% with 99% confidence. [4p
Answer #1

Solution :

Given that,

\hat p = 0.5

1 - \hat p = 1 - 0.5 = 0.5

margin of error = E = 2% = 0.02

At 99% confidence level the z is,

\alpha = 1 - 99%

\alpha = 1 - 0.99 = 0.01

\alpha/2 = 0.005

Z\alpha/2 = 2.576

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p)

= (2.576 / 0.02)2 * 0.5 * 0.5

=4147.36

Sample size = 4148